OpenStudy (anonymous):

These types of questions are hectic for me can someone help please |x - 7| - 7 = -3

OpenStudy (anonymous):

\[ |x-7|=4\\ x-7=\pm 4 \] Can you finish it?

OpenStudy (anonymous):

\[ x=7\pm 4\\ x=7+4=11\\ x=7-4=3 \]